# Improve Value Estimation of Q Function and Reshape Reward with Monte Carlo Tree Search

### v1

@@ Abstract
Reinforcement learning has achieved remarkable success in perfect information games such as Go and Atari, enabling agents to compete at the highest levels against human players. However, research in reinforcement learning for imperfect information games has been relatively limited due to the more complex game structures and randomness. Traditional methods face challenges in training and improving performance in imperfect information games due to issues like inaccurate Q value estimation and reward sparsity. In this paper, we focus on Uno, an imperfect information game, and aim to address these problems by reducing Q value overestimation and reshaping reward function. We propose a novel algorithm that utilizes Monte Carlo Tree Search to improve the value estimation in Q function. Even though we choose Double Deep Q Learning as the foundational framework in this paper, our method can be generalized and used in any algorithm which needs Q value estimation, such as the Actor-Critic. Additionally, we employ Monte Carlo Tree Search to reshape the reward structure in the game environment. We compared our algorithm with several traditional methods applied to games such as Double Deep Q Learning, Deep Monte Carlo and Neural Fictitious Self Play, and the experiments demonstrate that our algorithm consistently outperforms these approaches, especially as the number of players in Uno increases, indicating a higher level of difficulty.
@@ I Introduction
In previous research, there has been considerable exploration into the application of reinforcement learning in perfect information games, such as Go [14] [11], Chess[12] [13] and Atari[8] [20]. These studies have demonstrated significant success, with reinforcement learning agents consistently outperforming top human players through the deep neural networks to derive optimal policies. However, research in the domain of imperfect information games has been relatively scarce, and significant breakthroughs in this area remain limited. Several factors contribute to this difference. Imperfect information games are defined by the fact that players cannot access all the information within a given state, leading to varying levels of knowledge among participants, unlike in perfect information games. For instance, in most card games, players can only view their own hands, while crucial details, such as the cards held by opponents and those remaining in the deck, remain concealed. Additionally, these games are inherently non-deterministic and often include a high degree of randomness and luck, which presents significant challenges for reinforcement learning algorithms during training, potentially leading to convergence issues or high variance in results. Reinforcement learning agents in imperfect information games also encounter the challenge of sparse rewards, where obtaining rewards can be difficult, and positive feedback may be delayed for extended periods. This increases the computational demand for sampling and renders much of the sampled data ineffective. In previous research on imperfect information games like StarCraft[22][21], Doudizhu [27] [28] [29] and Texas Holdem [5] [1], Deep Monte Carlo(DMC)[18], Q-learning[24][20] and Neural Fictitious Self Play(NFSP)[5] have been the dominant and commonly used methods, yielding some results; however, they suffer from the issues such as overestimation of q values[20], inefficient for incomplete episodes and large variance of expected returns(rewards)[3], and often require an extensive number of training episodes to achieve satisfactory performance. These challenges highlight the necessity for new methods to improve convergence speed, reduce training steps and increase sample efficiency.
@@
    Play: If a player has a valid card, they must play it, setting it as the new target card, and the old target card is discarded. If they cannot play a valid card, they must draw a card from the draw pile. If the draw pile runs out, shuffle the discard pile to create a new draw pile.
@@
Reverse Reverse the direction of play; if it was initially clockwise, switch it to counterclockwise, and if it was counterclockwise, change it to clockwise.
@@
Figure 1: The different kind of Uno cards including the Number card, Skip, Reverse, Draw 2, Wild and Wild Draw 4
Uno, like other imperfect information games, suffers from the issue of reward sparsity, but to a more severe degree. According to the rules, only the player who plays all their cards first can win, which leads to situations where the deck is exhausted, requiring a reshuffle. This results in more rounds in Uno compared to other games(rounds shown in 2).
@@
In this paper, we aim to address the issues of reward sparsity and convergence challenges associated with problems of inaccurate estimation of Q values. To achieve this, we propose a new algorithm that improves value estimation of Q functions using Monte Carlo Tree Search (MCTS)[2] to solve Uno. Additionally, we utilize MCTS to reshape the reward structure within the Uno environment. We choose Double Deep Q-Learning (DDQN) as the base algorithm for training the agent and extend it to DDQN with MCTS; however, our method can be applied to any algorithm that requires Q value estimation. Our algorithm is trained for the same number of steps as DMC, NFSP, and DDQN in environments with three and four players. The evaluation results indicate that DDQN with MCTS achieves a higher win rate than the other algorithms by 4% - 16% with the same training episodes. More importantly, the integration of MCTS into DDQN exhibits a marked acceleration in performance improvement during the early and mid-training phases, whereas other algorithms may either show no significant progress or improve at a much slower rate.
@@ II-A Partially Observable Markov Decision Process
Any traditional reinforcement learning problems can be modelled as a mathematical framework called Markov Decision Process (MDP)[18], shown in Figure 3. In academic formalism, the MDP can be described as a five-tuple $(S,A,T,R,\gamma)$, where:
@@
  • $\gamma$ is the discount factor, which determines how future rewards are valued compared to immediate rewards, typically ranging from 00 to $1$.
Figure 3: The simplified Markov Decision Process (MDP) consists of four states, $S_{0}-S_{3}$, which represent the possible states. The actions, $a_{1}-a_{3}$, denote the decisions the agent makes to transition from one state to another. The transition function, $p$, represents the probability (greater than or equal to 0) of the agent taking a particular action at a state. The reward function, $r$, which can be either positive or negative, defines the reward the agent receives upon transitioning to a new state.
MDP assumes that all information about a state in the environment is accessible to agents. However, imperfect information games do not meet this condition, so they are modeled as a broader framework: Partially Observable Markov Decision Process (POMDP) [16]. A Partially Observable Markov Decision Process can be described as a seven-tuple $(S,A,T,R,\Omega,O,\gamma)$, where $(S,A,T,R,\gamma)$ are the same as in MDP, while the additional two elements can be defined as:
@@
  • $O(\omega|s^{\prime},a)$ is the observation probability function, which specifies the probability of observing $\omega$ when transitioning to state $s^{\prime}$ after taking action $a$.
The agent, based on the current state $s$ and observations $\Omega$, selects an action $a$, transitions to a new state $s^{\prime}$ according to the state transition probability $P$, receives a new observation based on $O(\omega|s^{\prime},a)$, and obtains a reward $R(s,a,s^{\prime})$. To solve POMDP, agents need to find an optimal policy $\pi^{*}$ that maximizes the expected cumulative reward over time. The optimal policy $\pi^{*}$ can be either deterministic or stochastic, selecting the best action given the current state and observations.
@@ II-B Q-learning and Deep Q-network
Q-learning[18] is one of the foundational and most widely-used algorithms in reinforcement learning, playing a critical role as both a core component and an essential building block for more advanced algorithms across various domains. One of the key characteristics of Q-learning is its reliance on the Bellman optimal equation for updates. With sufficient training data, Q-learning can guarantee that the agent converges to the optimal policy. As a model-free, bootstrapping algorithm, Q-learning learns to estimate the optimal action to take in any given state by calculating and updating Q-values—representing the expected utility of taking a particular action from a specific state. Q-learning operates on a Q-table where each state-action pair is associated with a Q-value. The algorithm updates its Q-values through the following steps in each iteration:
@@
    Update the Q-value: Update the Q-table using the following update rule:
    $Q(s,a)\leftarrow Q(s,a)+\alpha\left(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right)$
    where $s$ and $a$ are the original state and action taken, $s^{\prime}$ is the next state after taking the action, $r$ is the immediate reward received, $\alpha$ is the learning rate, and $\gamma$ is the discount factor, which determines the decay rate of the reward.
This iterative process allows the agent to learn an optimal strategy over time by refining its estimate of the Q-values for each state-action pair, ultimately converging on the policy that maximizes cumulative rewards.
Q-learning is well-suited for problems with small size of state and action spaces. However, in large-scale spaces, the traditional Q-learning method becomes impractical, as the Q-table can grow excessively large, making it inefficient to learn and store. To address this, Q-learning was replaced by Deep Q-Network(DQN) [9], which employs neural networks as Q value approximators to overcome the challenges of large state spaces. In addition, DQN introduced the technique of experience replay, where the agent’s experiences are stored in memory and used for batch updates rather than updating immediately after each transition. This approach increases sample efficiency by allowing the agent to learn from past experiences multiple times, improving the use of available data. Although DQN has become a benchmark algorithm in reinforcement learning, it has some drawbacks, one of the main issues being its tendency to overestimate Q-values. To address this, Double Deep Q-Network (DDQN) [20] was introduced. DDQN tackles the overestimation by using two separate neural networks: one network (estimator network) is used to predict the current Q-values for action selection, while the other network (target network), which is a past version of the estimator network, is used to calculate the target Q-values for updating. By separating these two processes and using a target network that lags behind the estimator, DDQN reduces the bias in estimating the target values, leading to more accurate and stable learning. It tries to minimize the loss between target value and estimated Q value to train the agent:
$L_{\text{ddqn}}=(Q_{\text{target}}(s,a;\theta^{\prime})-Q_{\text{estimator}}(s,a;\theta))^{2}$ (1)
@@
$Q_{\text{target}}(s,a;\theta^{\prime})$ $=r+\gamma Q_{\text{target}}\left(s^{\prime},\arg\max_{a^{\prime}}Q_{\text{estimator}}(s^{\prime},a^{\prime};\theta);\theta^{\prime}\right)$ (2)
@@ II-C Deep Monte Carlo and Monte Carlo Tree Search
    Initialize the return $R(s)$ of each state $s$ at time $t$ to 0. The average return for each $s$ is calculated using the formula $R(s_{t})=R(s_{t})+\lambda R(s_{t+1})$, where $\lambda$ is the reward discount factor.
@@
    Minimize the mean squared error between $R(s_{t})$ and $Q(s_{t},a_{t})$, where $Q(s_{t},a_{t})$ is predicted by the deep neural network. Repeat steps from 1 - 3 and finally, the optimal policy is generated by updating: For each $s$ in the episode, $\pi(s)\leftarrow\arg\max_{a}Q(s,a)$.
Monte Carlo Tree Search (MCTS) and DMC are both based on Monte Carlo sampling, but they are fundamentally different methods. MCTS builds a search tree by iterating over simulations of possible future moves and backpropagating the results, which makes it a key component of the renowned Go agent AlphaGo’s[11] algorithm. The MCTS process in reinforcement learning includes four principal stages:
@@
    Backpropagation: Once the simulation is complete, propagate the result back through the nodes along the path to the root. Update the visit counts and the average value estimates at each node based on the outcome of the simulation.
@@ III-A Uno Environment and Representations in Reinforcement Learning
We use RLCard [26], which is the Python game toolkit, as our Uno framework and reinforcement learning environment. One problem for our games is how to represent the game into a reinforcement learning environment. We have to define the elements of the MDP in the context of UNO:
@@
  • Reward (R): The reward is what a player gains after completing a round. In the basic rule of UNO, a reward of +1 is granted exclusively to the winner who has no cards left, while all others receive -1. During the ongoing game, the reward following all actions is 0. But our algorithm will reshape the reward structure based on MCTS, allowing the agent to receive rewards for certain actions taken.
@@ 60 Draw
We have adopted the state encoding approach described in RLCard, but our method differs by reducing the state size and omitting unnecessary information. While RLCard includes the player’s hand, the target card, and additional cards as part of the state, we argue that incorporating additional cards is redundant. Cards outside a player’s hand and target card include opponents’ cards and the deck, which is reshuffled in UNO. Since the agent can’t distinguish them with neural networks, this adds unnecessary complexity and may hinder training. Including additional cards would vastly expand the state space. To simplify and speed up training, we use only the player’s hand and target card. The state size of RLCard’s state encoding is $10^{126}$, while ours is $10^{72}$.
We encode hands and the target card into 4 planes. Every plane is a matrix of size 4x15, with each entry being 0 or 1 for one-hot encodings. The number 15 represents the different types of cards, disregarding color, which are the number cards from 0-9 and five kinds of special cards: skip, reverse, draw 2, wild, and wild 4. The number 4 represents the four colors: red, green, blue and yellow.
We encode the agent’s hand into three planes(shown in Figure 4), as in Uno, a player can have either 0, 1, or at most 2 cards of any given type. The plane 0 indicates that the player has zero of such cards in their hand. The plane 1 indicates that the player has exactly one of such cards. The plane 2 indicates that the player has two of such cards. For example, if the player has no Red 8 cards, the entry for Red 8 in the plane 0 is 1, while the entries for Red 8 in all other planes are 0. If the player has two Red 8 cards, the entry for Red 8 in the plane 2 is 1, and the entries for Red 8 in all other planes are 0. We encode the target card into a single plane where the matrix entry corresponding to the target card is 1, and the values of all other entries in the matrix are 0.

### v2

@@ Abstract
Reinforcement learning has achieved remarkable success in perfect information games such as Go and Atari, enabling agents to compete at the highest levels against human players. However, research in reinforcement learning for imperfect information games has been relatively limited due to the more complex game structures and randomness. Traditional methods face challenges in training and improving performance in imperfect information games due to issues like inaccurate Q value estimation and reward sparsity. In this paper, we focus on Uno, an imperfect information game, and aim to address these problems by reducing Q value overestimation and reshaping reward function. We propose a novel algorithm that utilizes Monte Carlo Tree Search to average the value estimations in Q function. Even though we choose Double Deep Q Learning as the foundational framework in this paper, our method can be generalized and used in any algorithm which needs Q value estimation, such as the Actor-Critic. Additionally, we employ Monte Carlo Tree Search to reshape the reward structure in the game environment. We compare our algorithm with several traditional methods applied to games such as Double Deep Q Learning, Deep Monte Carlo and Neural Fictitious Self Play, and the experiments demonstrate that our algorithm consistently outperforms these approaches, especially as the number of players in Uno increases, indicating a higher level of difficulty.
@@ I Introduction
In previous research, there has been considerable exploration into the application of reinforcement learning in perfect information games, such as Go [15] [12], Chess[13] [14] and Atari[9] [21]. These studies have demonstrated significant success, with reinforcement learning agents consistently outperforming top human players through the deep neural networks to derive optimal policies. However, research in the domain of imperfect information games has been relatively scarce, and significant breakthroughs in this area remain limited. Several factors contribute to this difference. Imperfect information games are defined by the fact that players cannot access all the information within a given state, leading to varying levels of knowledge among participants, unlike in perfect information games. For instance, in most card games, players can only view their own hands, while crucial details, such as the cards held by opponents and those remaining in the deck, remain unknown. Additionally, these games are inherently non-deterministic and often include a high degree of randomness and luck, which presents significant challenges for reinforcement learning algorithms during training, potentially leading to convergence issues or high variance in results. Reinforcement learning agents in imperfect information games also encounter the challenge of sparse rewards, where obtaining rewards can be difficult, and positive feedback may be delayed for extended periods. This increases the computational demand for sampling and renders much of the sampled data ineffective. In previous research on imperfect information games like StarCraft[23][22][7], Doudizhu [28] [30] [31] and Texas Holdem [5] [1], Deep Monte Carlo(DMC)[19], Q-learning[25][21] and Neural Fictitious Self Play(NFSP)[5][29] have been the dominant and commonly used methods, yielding some results; however, they suffer from the issues such as overestimation of Q values[21], inefficient for incomplete episodes and large variance of expected returns (rewards)[3], and often require an extensive number of training episodes to achieve satisfactory performance. These challenges highlight the necessity for new methods to improve estimation of values of a state, enhance convergence speed and increase sample efficiency.
@@
    Play: If a player has a valid card, they must play it, setting it as the new target card, and the old target card is discarded. If they cannot play a valid card, they must draw a card from the draw pile. If the draw pile runs out, shuffle the discarded cards to create a new draw pile.
@@
Reverse Reverse the direction of play: switch between clockwise and counterclockwise.
@@
Figure 1: The different kind of Uno cards including the Number card, Skip, Reverse, Draw 2, Wild and Wild Draw 4.
Uno, like other imperfect information games, suffers from the issue of reward sparsity, but to a more severe degree. The Uno rules lead to situations where the deck is exhausted, requiring a reshuffle, until one player plays all their cards and win. This results in more rounds in Uno compared to other games (round distribution shown in Figure 2).
@@
In this paper, we aim to address the issues of reward sparsity and convergence challenges associated with problems of inaccurate estimation of Q values. To achieve this, we propose a new algorithm that improves value estimation of Q functions using Monte Carlo Tree Search (MCTS)[2] to solve Uno. Additionally, we utilize MCTS to reshape the reward structure within the Uno environment. We choose Double Deep Q-Learning (DDQN) as the base algorithm for training the agent and extend it to DDQN with MCTS; however, our method can be applied to any algorithm that requires Q value estimation. Our algorithm is trained for the same number of steps as DMC, NFSP, and DDQN in environments with three and four players. The evaluation results indicate that DDQN with MCTS achieves a higher win rate than the other algorithms by 4% - 16%. More importantly, the integration of MCTS into DDQN exhibits a marked acceleration in performance improvement during the early and mid-training phases, whereas other algorithms may either show no significant progress or improve at a much slower rate.
@@ II-A Partially Observable Markov Decision Process
Any classic reinforcement learning problems can be modelled as a mathematical framework called Markov Decision Process (MDP)[19], shown in Figure 3. In academic formalism, the MDP can be described as a five-tuple $(S,A,T,R,\lambda)$, where:
@@
  • $\lambda$ is the discount factor, which determines how future rewards are valued compared to immediate rewards, typically ranging from 00 to $1$.
Figure 3: The simplified Markov Decision Process (MDP) consists of four states, $S_{0}-S_{3}$, which represent the possible states. The actions, $a_{1}-a_{3}$, denote the decisions the agent makes to transition from one state to another. The transition function, $p$, represents the probability (greater than or equal to 0) of the agent taking a particular action at a state. The reward function, $r$, which can be either positive or negative, defines the immediate reward the agent receives upon transitioning to a new state.
MDP assumes that all information about a state in the environment is accessible to agents. However, imperfect information games do not meet this condition, so they are modeled as a broader framework: Partially Observable Markov Decision Process (POMDP) [17]. POMDP can be described as a seven-tuple $(S,A,T,R,\Omega,O,\lambda)$, where $(S,A,T,R,\lambda)$ are the same as in MDP, while the additional two elements can be defined as:
@@
  • $O(\omega^{\prime}|s^{\prime},a)$ is the observation probability function, which specifies the probability of observing $\omega$ when transitioning to state $s^{\prime}$ after taking action $a$.
In POMDP, the agent, based on the current state $s$ and observations $\omega$, selects an action $a$, transitions to a new state $s^{\prime}$ according to the state transition probability $P$, receives a new observation based on $O(\omega^{\prime}|s^{\prime},a)$, and obtains a reward $R(s,a,s^{\prime})$. To solve POMDP, agents need to find an optimal policy $\pi^{*}$ that maximizes the expected cumulative reward over time. The optimal policy $\pi^{*}$ can be either deterministic or stochastic, selecting the best action given the current state and observations.
@@ II-B Q-learning and Deep Q-network
Q-learning[19] is one of the foundational and most widely-used algorithms in reinforcement learning, playing a critical role as both a core component and an essential building block for more advanced algorithms across various domains. One of the key characteristics of Q-learning is its reliance on the Bellman optimal equation for updates. With sufficient training data, Q-learning can guarantee that the agent converges to the optimal policy. As a model-free, bootstrapping algorithm, Q-learning learns to estimate the optimal action to take in any given state by calculating and updating Q values—representing the expected utility of taking a particular action from a specific state. Q-learning operates on a Q-table where each state-action pair is associated with a Q value. The algorithm updates its Q values through the following steps in each iteration:
@@
    Update the Q value: Update the Q-table using the following update rule:
    $Q(s,a)\leftarrow Q(s,a)+\alpha\left(r+\lambda\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right)$
    where $s$ and $a$ are the original state and action taken, $s^{\prime}$ is the next state after taking the action, $r$ is the immediate reward received, $\alpha$ is the learning rate, and $\lambda$ is the discount factor, which determines the decay rate of the future reward.
This iterative process allows the agent to learn an optimal strategy over time by refining its estimate of the Q values for each state-action pair, ultimately converging on the policy that maximizes cumulative rewards.
Q-learning is well-suited for problems with small size of state and action spaces. However, in large-scale spaces, the traditional Q-learning method becomes impractical, as the Q-table can grow excessively large, making it inefficient to learn and store. To address this, Q-learning was replaced by Deep Q-Network(DQN) [10], which employs neural networks as Q value approximators to overcome the challenges of large state and action spaces. In addition, DQN introduced the technique of experience replay, where the agent’s experiences are stored in memory and used for batch updates rather than updating immediately after each transition. This approach increases sample efficiency by allowing the agent to learn from past experiences multiple times, improving the use of available data. Although DQN has become a benchmark algorithm in reinforcement learning, it has some drawbacks, one of the main issues being its tendency to overestimate Q values. To address this, Double Deep Q-Network (DDQN) [21] was introduced. DDQN tackles the overestimation by using two separate neural networks: one network (estimator network) is used to predict the current Q values for action selection, while the other network (target network), which is a past version of the estimator network, is used to calculate the target Q values for updating. By separating these two processes and using a target network that lags behind the estimator, DDQN reduces the bias in estimating the target values, leading to more accurate and stable learning. It tries to minimize the loss between target value and estimated Q value to train the agent:
$L_{\text{ddqn}}=(Q_{\text{target}}(s,a;\theta^{\prime})-Q_{\text{estimator}}(s,a;\theta))^{2},$ (1)
@@
$Q_{\text{target}}(s,a;\theta^{\prime})$ $=r+\lambda Q_{\text{target}}\left(s^{\prime},\arg\max_{a^{\prime}}Q_{\text{estimator}}(s^{\prime},a^{\prime};\theta);\theta^{\prime}\right)$ (2)
@@ II-C Deep Monte Carlo and Monte Carlo Tree Search
    Initialize the return $G(s)$ of each state $s$ at time $t$ to 0. The average return for each $s$ is calculated using the formula $G(s_{t})=R(s_{t})+\lambda G(s_{t+1})$, where $\lambda$ is the reward discount factor.
@@
    Minimize the mean squared error between $G(s_{t})$ and $Q(s_{t},a_{t})$, where $Q(s_{t},a_{t})$ is predicted by the deep neural network. Repeat steps from 1 - 3 and finally, the optimal policy is generated by updating: For each $s$ in the episode, $\pi(s)\leftarrow\arg\max_{a}Q(s,a)$.
Another method based on Monte Carlo sampling is Monte Carlo Tree Search (MCTS), but unlike DMC, it builds a search tree by iterating over simulations of possible future moves and backpropagating the results, which makes it a key component of the renowned Go agent AlphaGo’s[12] algorithm. The MCTS process in reinforcement learning includes four principal stages:
@@
    Backpropagation: Once the simulation is complete, propagate the result back through the nodes along the path to the root. Update the visit counts at each node based on the outcome of the simulation.
@@ III-A Uno Environment and Representations in Reinforcement Learning
We use RLCard [27], which is the Python game toolkit, as our Uno framework and reinforcement learning environment. One problem for our method is how to represent the game into a reinforcement learning environment. We have to define the elements of the MDP in the context of UNO:
@@
  • Reward (R): The reward is what a player gains after completing a round. In the basic rule of UNO, a reward of +1 is granted exclusively to the winner who has no cards left, while all others receive -1 at the end. During the ongoing game, the reward following all actions is 0. But our algorithm will reshape the reward structure based on MCTS, allowing the agent to receive rewards for certain actions taken.
@@ 60 Draw
We have adopted the state encoding approach described in RLCard, but our method differs by reducing the state size and omitting unnecessary information. While RLCard includes the player’s hand, the target card, and additional cards as part of the state, we argue that incorporating additional cards is redundant. Cards outside a player’s hand and target card include opponents’ cards and the deck, which is reshuffled in UNO. Since the agent can’t distinguish them in neural networks, this adds unnecessary complexity and may hinder training. Including additional cards would also vastly expand the state space. To simplify and speed up training, we use only the player’s hand and target card. The state size of RLCard’s state encoding is $10^{126}$, while ours is $10^{72}$.
We encode hands and the target card into 4 planes. Every plane is a matrix of size 4x15, with each entry being 0 or 1 for one-hot encodings. The number 15 represents the different types of cards, disregarding color, which are the number cards from 0-9 and five kinds of special cards: skip, reverse, draw 2, wild, and wild 4. The number 4 represents the four colors: yellow, green, blue and red.
The agent’s hand information is encoded into three planes(shown in Figure 4), as in Uno, a player can have either 0, 1, or at most 2 cards of any given type. The plane 0 indicates that the player has zero of such cards in their hand. The plane 1 indicates that the player has exactly one of such cards. The plane 2 indicates that the player has two of such cards. For example, if the player has no Red 8 cards, the entry for Red 8, located in the fourth row and the ninth column, in the plane 0 is 1, while the entries for Red 8 in all other planes are 0. If the player has two Red 8 cards, the entry for Red 8 in the plane 2 is 1, and the entries for Red 8 in all other planes are 0. We encode the target card into a single plane where the matrix entry corresponding to the target card is 1, and the values of all other entries in the matrix are 0.